PROBEUSERS Archives

June 2011

PROBEUSERS@LISTS.UMN.EDU

Options: Use Monospaced Font
Show HTML Part by Default
Show All Mail Headers

Message: [<< First] [< Prev] [Next >] [Last >>]
Topic: [<< First] [< Prev] [Next >] [Last >>]
Author: [<< First] [< Prev] [Next >] [Last >>]

Print Reply
Subject:
From:
Joe Geller <[log in to unmask]>
Reply To:
JEOL-Focused Probe Users List <[log in to unmask]>
Date:
Tue, 14 Jun 2011 09:56:38 -0400
Content-Type:
multipart/alternative
Parts/Attachments:
text/plain (3262 bytes) , text/html (7 kB)
JEOL Probe Users Listserver

Moderator: Ellery Frahm, [log in to unmask],
Electron Microprobe Lab, University of Minnesota

Post a message: send your message to [log in to unmask]

Unsubscribe: send "SIGNOFF PROBEUSERS" to [log in to unmask]

On-line help and FAQ: http://probelab.geo.umn.edu/listserver.html



*

Jim:

                Check the filament continuity. It is best done by removing
the HV cable from the oil tank (make sure the power is off and ground the
connectors) so the cable, contacts and filament are all tested together.

 

 

Joe Geller

Geller MicroAnalytical Lab., Inc.

426e Boston St., Topsfield, MA 01983

www.GellerMicro.com

978 887-7000

 

 

 

From: JEOL-Focused Probe Users List [mailto:[log in to unmask]] On
Behalf Of Jim Renaud
Sent: Monday, June 13, 2011 7:19 PM
To: [log in to unmask]
Subject: [PROBE-USERS] JEOL 733 issue

 

JEOL Probe Users Listserver 

Moderator: Ellery Frahm, [log in to unmask], Electron Microprobe Lab,
University of Minnesota 

Post a message: send your message to [log in to unmask] 

Unsubscribe: send "SIGNOFF PROBEUSERS" to [log in to unmask] 

On-line help and FAQ: http://probelab.geo.umn.edu/listserver.html 

* 

Hi Everyone,

 

    I am running a JEOL 733 and have run into some difficulties.  Last
tuesday I did a sample exchange before closing the office up for the night.
A couple of hours later I realized that there was no current on the cup, so
I figured it was time for a filament change.  I inspected the old filament
and realized it had not blown at all.  I installed a new filament and pumped
the probe down.  The probe pumped down, 15KV was achieved.  Normally when I
have run the probe, the coarse gain knob on the condenser lens has been on
step 3 or 4,  but now no cup current or absorbed current are achieved at
this setting.  I adjusted the aperature with no significant change.  I
cycled through the coarse gain knob of the condenser lens (on the Electron
Optical System (EOS)) from 1-7 and at step 8 (one step above 7 on the knob
but not labelled),  I obtained picoamps of current on the cup.  At one
point, i did achieve 12 nano-amps at the step 7 setting, but that
disappeared.  Since I was not able to  get any current on any of the other
steps while cycling through the condenser knob, I figured there was a
problem with the condenser lens.  I swapped the EOS module, the ML cooling
Box at the bottom of the console, the actual condenser lens, and the
accelerating voltage module (thank goodness for spare parts), however, no
change in the current, and again, I could only get picoamps of current on
step 7 or 8.  We have checked voltages down at the ML cooling box, and at
various other locations and they seem to coincide with the JEOL circuit
diagrams.  We then decided to investigate the brass/copper tube liners which
run the center of the column, cleaned them and re-installed them (are these
column tube liners supposed to be grounded?).  Again no change in the
current on the cup nor on the sample and only picoamps achieved at step 8 on
the condenser lens.  I am running out of things to try.  Some other
possibilities I am considering is the potential of charging somewhere in the
column; perhaps the leaded OM glass has become charged?  Any thoughts,
suggestions, or things to try would be much appreciated.

 

Thanks,

 

Jim.



ATOM RSS1 RSS2